### v2

# AI4Science: The Reach and Limits of AI for Scientific Discovery

Abstract
Simulating the long-timescale dynamics of biomolecules is a central challenge in computational science. While enhanced sampling methods can accelerate these simulations, they rely on pre-defined collective variables that are often difficult to identify. A recent generative model, LD-FPG, demonstrated that this problem could be bypassed by learning to sample the static equilibrium ensemble as all-atom deformations from a reference structure, establishing a powerful method for all-atom ensemble generation. However, while this approach successfully captures a system’s probable conformations, it does not model the temporal evolution between them. Here we extend LD-FPG with a temporal propagator that operates within the learned latent space and compare three classes: (i) score-guided Langevin dynamics, (ii) Koopman-based linear operators, and (iii) autoregressive neural networks. Within a unified encoder–propagator–decoder framework, we evaluate long-horizon stability, backbone and side-chain ensemble fidelity, and functional free-energy landscapes. Autoregressive neural networks deliver the most robust long rollouts; score-guided Langevin best recovers side-chain thermodynamics when the score is well learned; and Koopman provides an interpretable, lightweight baseline that tends to damp fluctuations. These results clarify the trade-offs among propagators and offer practical guidance for latent-space simulators of all-atom protein dynamics.
1 Introduction
Molecular simulations are indispensable for studying the complex dynamics that govern biological function, yet brute-force approaches struggle to access the slow, functionally relevant motions—such as protein folding, ligand binding, or allosteric switching—due to rugged energy landscapes and the dominance of rare events (Dror et al., 2011; Latorraca et al., 2017). To mitigate this gap—beyond what enhanced sampling can offer when suitable collective variables are hard to specify—a complementary strategy has gained traction: shifting the burden from raw coordinates to learned latent coordinates. In this representation-first view, the simulation problem is recast as a modular encoder–propagator–decoder pipeline: an encoder maps high-dimensional atomic configurations into a continuous, low-dimensional latent space; a propagator evolves the system’s state within this simplified space; and a decoder maps the resulting latent trajectory back to all-atom coordinates (Sidky et al., 2020; Wu et al., 2018).
Progress within this paradigm comes from two complementary directions. The first focuses on learning the underlying physics, employing score-based diffusion, flow matching, and energy-based models to learn generative surrogates—sometimes yielding differentiable force fields—that implicitly define the system’s potential of mean force (Arts et al., 2023; Plainer et al., 2025). The second centers on learning simplified dynamical coordinates, using time-aware autoencoders or Koopman/DMD analysis to discover an intrinsic manifold where the long-term dynamics become stable, predictable, or even approximately linear (Hernández et al., 2018; Mardt et al., 2018; Brunton et al., 2021; Azencot et al., 2020).
While this framework is well-established, the choice of the latent propagator—the engine that drives the dynamics—remains a critical open question, as different models offer distinct trade-offs between physical rigor, long-term stability, and expressive power. Within this paradigm, we adopt LD-FPG (Sengar et al., 2025) as the encoder–decoder backbone that learns an all-atom equilibrium ensemble in a pooled latent space, and we augment it with a temporal component. We then systematically compare, within the same latent space, three propagator classes: (i) score-guided Langevin dynamics, which leverages learned forces from the equilibrium distribution; (ii) Koopman-based linear operators, which offer long-horizon stability and interpretability; and (iii) flexible neural networks (MLPs), which capture non-linear memory effects but can drift during long autoregressive rollouts (Vlachas et al., 2021; Greydanus et al., 2019; Cranmer et al., 2020).
We evaluate these propagators on alanine-dipeptide and two GPCRs (A1AR, A2AR), assessing long-horizon stability, backbone and side-chain ensemble fidelity, and functional free-energy landscapes. In brief, autoregressive neural networks provide the most reliable long rollouts; score-guided Langevin best recovers side-chain thermodynamics when the score is well learned; and Koopman serves as a lightweight, interpretable baseline that tends to damp fluctuations.
2 Related Work
The encoder–propagator–decoder blueprint.
The core idea of simulating in a low-dimensional space is well established. Molecular Latent Space Simulators (LSS) (Sidky et al., 2020) explicitly factorize the problem into three components: an encoder to find slow collective variables (CVs), a latent propagator to evolve them, and a decoder to generate all-atom structures. Similarly, Deep Generative MSMs (DeepGenMSM) (Wu et al., 2018) pair a latent Markovian transition model with a generative decoder to emit molecular configurations for each state. This blueprint has been realized in various forms, including autoregressive simulators with RNN/LSTM propagators (Vlachas et al., 2021), trajectory-level generators that cast MD as video synthesis (Jing et al., 2024), and invertible models like Boltzmann Generators that learn a direct map to the equilibrium distribution (Noé et al., 2019). We follow this modular design but (i) hold the LD-FPG decoder fixed to control for reconstruction quality and (ii) focus our analysis on the choice of latent propagator.
Learning dynamically aware latent spaces (encoders).
The quality of a latent simulation hinges on the quality of the latent space itself. While linear methods like time-lagged independent component analysis (TICA) remain strong baselines for identifying slow variables (Pérez-Hernández et al., 2013), deep learning has enabled far more expressive encoders. Time-lagged autoencoders such as the Variational Dynamics Encoder (VDE) learn non-linear representations predictive over a time delay $\Delta t$ (Hernández et al., 2018; Wehmeyer and Noé, 2018). VAMPnets use a variational principle to approximate the leading Koopman eigenfunctions, corresponding to the slowest dynamical processes (Mardt et al., 2018). Similarly, iterative methods like RAVE use VAEs to learn a latent distribution that guides the discovery of an optimal reaction coordinate for enhanced sampling (Ribeiro et al., 2018). To remove hand-crafted features, modern approaches leverage graph neural networks (GNNs) to learn CVs directly from coordinates in a permutation- and symmetry-aware manner (Zou et al., 2025; Ghorbani et al., 2022; Zhang et al., 2024; Duval et al., 2023), often with information-bottleneck objectives to explicitly optimize predictiveness of future states (Wang and Tiwary, 2021).
Linear dynamics in latent space (Koopman/DMD propagators).
The Koopman-operator framework provides a route to linearize non-linear dynamics in a learned observable space (Brunton et al., 2021). Extended dynamic mode decomposition (EDMD) (Williams et al., 2015) and DMD (Tu, 2013) showed that a wide class of systems admits accurate linear predictors; subsequent work learns such observables end-to-end with deep encoders (Lusch et al., 2018). This has inspired Koopman autoencoders that enforce a linear evolution rule within the latent space (Tayal et al., 2023; Wu et al., 2025), yielding exceptional long-horizon stability by avoiding the compounding errors of iterated non-linear models, with continued advances in scalable kernels and consistency-enforcing architectures (Azencot et al., 2020).
Non-linear sequence models (neural propagators).
A complementary path directly learns the non-linear transition function. The Learning Effective Dynamics (LED) framework uses LSTMs to propagate latent variables and capture memory effects (Vlachas et al., 2021); others pair RNNs with physical resampling such as Maximum Caliber to enforce kinetic consistency (Tsai et al., 2022). Continuous-time learners (Neural ODEs) offer flexible parametric flows (Chen et al., 2018), and graph-based simulators exploit locality for rollouts in interacting systems (Sanchez-Gonzalez et al., 2020). To curb instability, physics-preserving architectures (Hamiltonian/Lagrangian networks) enforce conservation laws (Greydanus et al., 2019; Cranmer et al., 2020). Our MLP propagator serves as a streamlined baseline in this family.
Physics-guided stochastic models (Langevin and diffusion).
Generative models increasingly embed statistical-mechanics structure. Score-based diffusion has been used to learn effective force fields for coarse-grained MD (Arts et al., 2023), linking denoising, Langevin dynamics  (Arts et al., 2023), and Fokker–Planck evolution (Plainer et al., 2025). Recent works (e.g., DiffMD, Score Dynamics) demonstrate larger stable time steps while retaining short-time kinetic signals (Wu and Li, 2023; Hsu et al., 2024). A key theme is consistency between the sampled equilibrium ensemble and the stationary distribution of the learned dynamics (Plainer et al., 2025); stability can also benefit from noise augmentation during training (Arts et al., 2023).
GPCRs as a proving ground for slow dynamics.
G protein-coupled receptors (GPCRs) are a demanding testbed: their function hinges on slow transitions among metastable states mapped by landmark MD/MSM studies (Dror et al., 2011; Kohlhoff et al., 2014; Latorraca et al., 2017). High-resolution structures (e.g., $\beta_{2}$AR–Gs) provide anchors for validating pathways (Rasmussen et al., 2011), and reviews emphasize micro-switches and long-range allostery spanning orders of magnitude in time (Miao and McCammon, 2016; Manglik et al., 2015; Venkatakrishnan et al., 2016). By evaluating on A1AR/A2AR, we benchmark whether latent simulators recover metastable states and kinetic pathways relevant to activation.
Positioning among recent generative simulators.
Recent efforts span coordinate-space trajectory synthesis (e.g., diffusion-based GeoTDM) (Han et al., 2024), SE(3)-equivariant flow matching for coarse-grained rollouts (F3low) (Li et al., 2024), and neural operators for full 3D dynamics beyond next-step prediction (EGNO) (Xu et al., 2024). Latent and physics-informed approaches accelerate sampling or enforce structure—LAST for adaptive MD (Tian et al., 2022), ConfRover for joint conformation–dynamics learning (Shen et al., 2025), and NeuralMD with symmetric neural ODEs for binding dynamics (Liu et al., 2024)—while comparative studies benchmark diffusion, flow matching, and normalizing flows on MD tasks (John et al., 2025). Complementing these advances, this work isolates the propagator choice by evaluating linear (Koopman), neural, and score-guided dynamics within the same learned latent and fixed decoder, enabling a controlled comparison of long-horizon stability, ensemble fidelity, and functional landscapes. We also considered MD-Gen (Jing et al., 2024), but as it was pre-trained on the Atlas dataset, which does not include alanine-dipeptide, re-training it for a fair comparison was beyond the scope of this work.
3 Methods
Figure 1: Framework overview. A pre-trained LD-FPG encoder (ChebNet; left) maps all-atom coordinates $X(t)$ to a pooled latent $z(t)$. Within this fixed latent, one of three propagators advances the state (red box): (a) score-guided Langevin using the LD-FPG denoiser to estimate $s_{\theta}(z,\tau)=\nabla_{z}\log p_{\tau}(z)$ at a fixed low-noise level; (b) an autoregressive neural network $z_{t+1}=f_{\theta}(z_{t})$; and (c) a Koopman linear operator $z_{t+1}=Az_{t}$. The frozen LD-FPG decoder (right) maps the latent trajectory back to all-atom structures $\hat{X}(t+\Delta t)$
Our framework extends the Latent Diffusion for Full Protein Generation (LD-FPG) model by incorporating a temporal propagator that operates within its learned conformational latent space. To provide a self-contained description, we first briefly outline the relevant components of the LD-FPG architecture before detailing the three propagator models developed in this work.
3.1 Latent Space Representation from LD-FPG
The foundation of our method is the encoder-decoder architecture from LD-FPG. An encoder, implemented as a Chebyshev Graph Neural Network (ChebNet), learns a mapping from high-dimensional, all-atom protein coordinates $X(t)\in\mathbb{R}^{N\times 3}$ (where NN is the number of heavy atoms) to a low-dimensional latent embedding $z(t)\in\mathbb{R}^{d}$. This encoder is trained on a Molecular Dynamics (MD) trajectory, producing a time-series of latent vectors $\{z_{0},z_{1},\dots,z_{M}\}$ that captures the essential conformational dynamics of the protein. Our goal is to model the time evolution within this latent space, which can be expressed as a discrete-time update rule:
$z_{t+1}=f(z_{t})+\eta_{t}$ (1)
where $f$ is the propagator function we aim to learn, and $\eta_{t}$ represents a stochastic noise term. We systematically compare three distinct classes of propagators for learning $f$. Key notation is summarized in Table S1, and a glossary in Table S2.
We use a one–frame latent stride for training pairs $(z_{t},z_{t+1})$. Unless stated, rollout noise is $\eta_{t}\sim\mathcal{N}(0,I)$.
3.2 Koopman Propagator via Dynamic Mode Decomposition
We approximate the latent dynamics with a linear map
$z_{t+1}\approx Az_{t},\qquad A\in\mathbb{R}^{d\times d}.$ (2)
Let the snapshot matrices collect columns as time:
$\mathbf{X}=[\,z_{0},\dots,z_{M-2}\,]\in\mathbb{R}^{d\times(M-1)},\quad\mathbf{Y}=[\,z_{1},\dots,z_{M-1}\,]\in\mathbb{R}^{d\times(M-1)}.$
DMD solves $\min_{A}\|\mathbf{Y}-A\mathbf{X}\|_{F}^{2}$ with closed form
$A=\mathbf{Y}\,\mathbf{X}^{+},$ (3)
where + is the Moore–Penrose pseudoinverse. For stability, we compute $\mathbf{X}^{+}$ via truncated SVD of $\mathbf{X}$ at rank $r<d$ (EDMD); $r$ is chosen by retaining singular values above a fixed energy fraction (e.g., $95\%$). New trajectories follow $\hat{z}_{t+1}=A\hat{z}_{t}+\eta_{t}$ with optional $\eta_{t}\sim\mathcal{N}(0,\sigma^{2}I)$.
3.3 Autoregressive Neural Network Propagator
To capture potentially complex, non-linear relationships in the dynamics, we employ a standard Multi-Layer Perceptron (MLP) as the propagator. The model learns a general non-linear function $f_{\theta}$ parameterized by weights $\theta$:
$z_{t+1}=f_{\theta}(z_{t})$ (4)
Our implementation of $f_{\theta}$ is a sequential network consisting of fully-connected layers with ReLU activation functions and Dropout for regularization. The model is trained to predict the state one step ahead by minimizing the Mean Squared Error (MSE) loss between the predicted state and the true next state over the training data:
$\mathcal{L}(\theta)=\frac{1}{M-1}\sum_{t=0}^{M-2}\|f_{\theta}(z_{t})-z_{t+1}\|^{2}$ (5)
The optimization is performed using the Adam optimizer. Similar to the Koopman model, trajectories are generated autoregressively from a starting point $\hat{z}_{0}$ by iteratively applying the learned function: $\hat{z}_{i+1}=f_{\theta}(\hat{z}_{i})+\eta_{i}$.
3.4 Score-Guided Langevin Propagator
This approach frames the dynamics from a statistical mechanics perspective, simulating the evolution of the system under the influence of a potential of mean force. The dynamics are governed by the overdamped Langevin equation, which in discretized form is:
$z_{t+1}=z_{t}-\nabla_{z}U(z_{t})\,h+\sqrt{2Th}\,\eta_{t},$ (6)
where $U(z)$ is the potential of mean force, $\Delta t$ is the integration time step, $T$ is the temperature, and $\eta_{t}\sim\mathcal{N}(0,I)$ is a random Gaussian vector.
The key insight is that the force term, $\nabla_{z}U(z)$, can be related to the score of the equilibrium (Boltzmann) distribution, $s(z)=\nabla_{z}\log p(z)$, since $p(z)\propto\exp(-U(z)/T)$. The pre-trained diffusion model from LD-FPG, $\epsilon_{\theta}(z_{\tau},\tau)$, provides a direct way to estimate this score. According to score-based generative modeling theory, the score of a data distribution perturbed with noise level $\sigma_{\tau}$ is related to the optimal denoiser:
$\nabla_{z}\log p_{\tau}(z_{\tau})\approx-\frac{\epsilon_{\theta}(z_{\tau},\tau)}{\sigma_{\tau}^{2}}.$ (7)
By evaluating the model at a low, fixed noise level (i.e., a small diffusion timestep $\tau_{\text{noise}}$), we approximate the score of the true data distribution, $s(z)\approx-\frac{\epsilon_{\theta}(z,\tau_{\text{noise}})}{\sigma_{\tau_{\text{noise}}}^{2}}$ (song2021score; De Bortoli et al., 2021).
Substituting the score for the force term, we arrive at the simulation update rule:
$z_{i+1}=z_{i}+Th\,s(z_{i})+\sqrt{2Th}\,\eta_{i},\qquad\eta_{i}\sim\mathcal{N}(0,I).$ (8)
This method directly leverages the learned equilibrium distribution from LD-FPG to drive a physically-motivated, stochastic simulation in the latent space. To enhance numerical stability during long rollouts, we also implement optional score clipping, where the norm of the score vector $s(z_{i})$ is capped at a predefined maximum value.
Figure 2: Stability over long rollouts. RMSD and lDDT versus frame index for (a) alanine-dipeptide and (b) A1AR. The failure time is defined as the first frame where lDDT (computed relative to the first frame of the trajectory) drops below 0.65. Autoregressive NN maintains the longest stable rollout on A1AR (no failure within 10,000 frames), while Koopman and Langevin fail earlier; on alanine-dipeptide, Koopman and NN persist for thousands of frames whereas Langevin fails early.
4 Results and Discussion
We benchmark the three latent-space propagators—Koopman, Autoregressive Neural Network (NN), and score-guided Langevin—within the unified LD-FPG latent (Fig. 1). Datasets and code are summarized in Appendix S1.2. Metrics include stability (RMSD, lDDT), equilibrium-ensemble fidelity (backbone and side-chain dihedral JSD), and functional free-energy landscapes. Unless noted, failure time is the first frame where lDDT relative to the initial frame drops below 0.65.
4.1 Long-horizon stability: autoregressive NN is most robust
Figure 2 tracks RMSD and lDDT through time; failure time is the first sampled point where lDDT (vs. the first frame) drops below $0.65$. For alanine-dipeptide (Fig. 2a), Koopman and the Autoregressive NN remain stable for several thousand sampled frames (failure at 4443 and 3176, respectively), whereas Langevin fails early (206). On A1AR (Fig. 2b), the Autoregressive NN completes the full 10,000-frame rollout without failure; Langevin remains stable to 7476; Koopman fails at 5740. A2AR (Table 1; SI) follows the same ranking: Autoregressive NN >> Langevin >> Koopman for long-rollout stability.
A note on “frames” for Langevin.
For Koopman and NN, each frame corresponds to the dataset stride (the models are trained to map $z_{t}\!\to\!z_{t+1}$), so the horizontal axis coincides with the MD sampling stride. By contrast, the Langevin propagator integrates a latent SDE with internal step size $\Delta t$ and a separate sampling stride; thus a “frame” is a sampled SDE state, not a single update of the integrator. To make curves visually comparable, we calibrated $(\Delta t,\text{sampling stride})$ per system to match the short-horizon RMSD across Langevin replicas to the base simulation. Concretely, for alanine-dipeptide we used $\Delta t=10^{-8}$ and sampled every step (one internal step per plotted frame), whereas for A1AR we used $\Delta t=10^{-10}$ and sampled every 20 steps (effective time per plotted frame $20\,\Delta t$). The reported failure indices for Langevin therefore count sampled outputs; converting to physical time would scale the A1AR axis by $20\,\Delta t$ and the alanine axis by $\Delta t$.
Why Langevin fails early on alanine.
The rapid failure of Langevin dynamics on the dipeptide likely stems from a combination of challenges in the training data, the system’s intrinsic properties, and the simulation hyperparameters. First, the large time step between frames in the source MD simulation can result in a sparsely sampled, fragmented latent manifold. This makes it difficult for the diffusion model to learn a smooth and continuous score function ($s(z)=\nabla_{z}\log p(z)$), yielding a noisy or inaccurate effective force field that is prone to instability. Second, the dipeptide’s small size leads to large intrinsic fluctuations and high successive-frame displacements (RMSD $\sim\!0.9$ Å), making the lDDT failure criterion particularly stringent. Finally, the simulation hyperparameters were reused from GPCR settings (Sengar et al., 2025). The large effective step size, chosen to model such highly diffusive systems, proves too aggressive when combined with the imperfect score function. This combination rapidly drives the simulation into unphysical regions, a deviation that is quickly detected by the sensitive lDDT metric.
4.2 Equilibrium-ensemble fidelity: NN best on backbone; Langevin best on side-chains
We evaluate how well each propagator reproduces the equilibrium ensemble in dihedral space using 2D free-energy maps for backbone $(\phi,\psi)$ and side-chain $(\chi_{1},\chi_{2})$ angles (Fig. 3). Fidelity is quantified with the Jensen–Shannon divergence (JSD) between model and ground-truth distributions (Table 1); side-chain JSD is aggregated over all residues so that it captures global rotamer statistics.
Alanine-dipeptide.
All three models recover the canonical Ramachandran basins. The Autoregressive NN most closely matches basin shape and separation (backbone JSD =0.0056=0.0056), slightly outperforming Koopman ($0.0085$). Langevin under-samples and over-smooths the landscape (backbone JSD =0.029=0.029), consistent with its early rollout failure on this small, rapidly fluctuating system (Fig. 2a). For comparison, an Equivariant Graph Neural Operator (EGNO) baseline (Xu et al., 2024) yielded a significantly higher backbone JSD of $0.3875$, underscoring the effectiveness of our specialized latent-space propagators.
A1AR
: For the large GPCR, the Autoregressive NN gives the most faithful backbone ensemble (backbone JSD =0.0443=0.0443), while Langevin broadens low-energy regions (backbone JSD =0.1943=0.1943). In contrast, side-chains tell the opposite story: the score-guided Langevin dynamics sharply recovers rotameric structure in $(\chi_{1},\chi_{2})$ with the lowest divergence (JSD =0.0223=0.0223), the NN is second-best ($0.0436$), and Koopman is notably diffuse ($0.1144$). Visually, the Langevin maps reproduce the expected $\chi_{1}$ bands (g+, anti, g-) present in the ground truth, whereas the NN blurs band edges and Koopman largely washes them out (Fig. 3, right).
Why side-chains favor Langevin.
Side-chain rotamers are governed by local barriers and short-range couplings. Because the Langevin propagator uses the learned score $s(z)=\nabla_{z}\log p(z)$ from LD-FPG, it injects a thermodynamically consistent drift toward high-probability neighborhoods in latent space and adds small isotropic noise. This combination encourages frequent, barrier-crossing micro-moves that preserve the stationary rotamer distribution. The NN, optimized for one-step prediction, can accumulate exposure bias and smooth sharp multimodality over long rollouts; Koopman’s linear evolution further damps variance, leading to broadened or merged rotamer basins.
Figure 3: Ensemble fidelity in dihedral space. Free-energy maps for backbone $(\phi,\psi)$ and, for A1AR, side-chain $(\chi_{1},\chi_{2})$ angles. Alanine-dipeptide (left) shows recovery of the canonical Ramachandran basins by all methods, with the Autoregressive NN most closely matching basin shapes given the long rollout. For A1AR (middle/right), NN best matches backbone statistics, while score-guided Langevin achieves the lowest divergence for side-chains (side-chain JSD aggregated over all residues).
Amplitude of motion.
Across systems, Koopman systematically underestimates fluctuations (lower mean RMSF than ground truth), reflecting its variance-damping bias, whereas Langevin and the NN better match the amplitude of motion (Table 1). Residue-resolved trends follow the same pattern (see Fig. S1).
4.3 Functional GPCR surfaces: TM-distance free energies capture activation
[…]
A1AR.
Both the NN and Langevin models recover the location and curvature of the principal low–free-energy valley. Langevin spans the valley most extensively, covering the transition corridor between inactive and active-like states. The NN tracks the same valley with a tighter footprint, under-sampling the flanks. Koopman identifies the basin center but exhibits stiffer, more isotropic contours and elevated apparent barriers, limiting coverage of the transition path. The one-dimensional slices mirror these trends: NN and Langevin reproduce the primary minimum and overall shape, whereas Koopman inflates barriers.
A2AR (reference).
For A2AR (Fig. S2), switching involves coordinated changes along both axes: TM3–6 increases with the TM6 outward motion and TM3–7 shifts as the intracellular pocket reshapes, yielding a diagonal valley in the ($\mathrm{TM3\text{--}6}$, $\mathrm{TM3\text{--}7}$) plane. Langevin again achieves the broadest coverage along both coordinates and reaches the transition corridor; the NN follows the valley with narrower support; Koopman under-covers the corridor and smooths anisotropy.
Summary.
Together with the dihedral analyses, these surfaces indicate that the non-linear (NN) and score-guided (Langevin) propagators better explore and connect metastable states relevant to GPCR activation, while the strictly linear Koopman rule provides a conservative baseline that tends to over-regularize barriers and shrink anisotropy.
Figure 4: Functional free-energy surface for A1AR. Two-dimensional free-energy over TM3–6 (“TM36”) and TM3–7 (“TM37”) distances. The background heat map shows the ground-truth surface; overlaid contours show Koopman (purple), Autoregressive NN (cyan), and Langevin (orange). Right panels: corresponding 1D free-energy profiles along each coordinate. NN and Langevin track the main basin and curvature; Koopman identifies the basin but overestimates barriers and misses anisotropy.
Table 1: Quantitative summary across systems. Failure time is the first frame with lDDT (vs. the initial frame) <0.65<0.65. JSDs measure divergence from ground-truth dihedral distributions. Average RMSF is in Å. For alanine-dipeptide, we include an EGNO baseline (Xu et al., 2024) trained for 100 frames. A2AR results are in the SI.
System / Metric Model Failure Time (frames) Backbone JSD Sidechain JSD Average RMSF (Å)
Alanine-Dipeptide Ground Truth N/A N/A N/A 0.8222
Koopman 4443 0.0085 N/A 0.8320
Autoregressive NN 3176 0.0056 N/A 0.8171
Langevin Dynamics 206 0.029 N/A 0.8233
EGNO (baseline) 1000 0.3875 N/A 1.0481
A1AR Ground Truth N/A N/A N/A 1.5809
Koopman 5740 0.0472 0.1144 0.7482
Autoregressive NN No failure (10000) 0.0443 0.0436 0.7880
Langevin Dynamics 7476 0.1943 0.0223 1.1303
A2AR Ground Truth N/A N/A N/A 1.9454
Koopman 2324 0.1195 0.1380 1.2531
Autoregressive NN 5789 0.0679 0.0691 0.8200
Langevin Dynamics 5432 0.1211 0.0065 0.7463
4.4 Outlook
Across systems, the propagators fill complementary roles. The autoregressive NN provides the most reliable long-horizon rollouts and the closest backbone statistics (it is the only A1AR run without failure). Score-guided Langevin best reproduces side-chain thermodynamics on GPCRs and yields realistic local fluctuations, though it is sensitive to score quality and step size. Koopman remains a fast, interpretable baseline that offers medium-term stability but damps variance and blurs rotamer structure. In practice, we favor the NN when long all-atom trajectories are the goal and Langevin when rotamer distributions and local thermodynamics are paramount, provided the denoiser and latent connectivity are well tuned.
Appendix S1 Supplementary Information
S1.1 Supplementary Figures for A2AR
Figure S1: A2AR dihedral statistics (SI). Backbone $(\phi,\psi)$ and side-chain $(\chi_{1},\chi_{2})$ free-energy maps for A2AR. Trends mirror the main text: Autoregressive NN provides the best backbone match among learning-based models, while score-guided Langevin attains the lowest divergence for side-chains (side-chain JSD aggregated over all residues).
Figure S2: A2AR TM-distance free-energy (SI). Two-dimensional free-energy over TM3–6 (“TM36”) and TM3–7 (“TM37”) distances. The background is the ground-truth surface; three overlapping contour sets show Koopman (purple), Autoregressive NN (cyan), and Langevin (orange). Right panels: 1D free-energy profiles along each coordinate. Consistent with A1AR, NN and Langevin follow the principal basin and curvature more closely than Koopman.
S1.2 Simulation Datasets and Code availability
The complete code for the LD-FPG framework is publicly available on GitHub: https://github.com/adityasengar/LD-FPG/ and the code for the latent propagator is available here: https://github.com/adityasengar/latent-dynamics-propagators
The latent trajectories used to train the propagators were derived from three publicly available Molecular Dynamics (MD) simulation datasets. For each system, the original source provided a structure file (PDB) and a coordinate trajectory (XTC). These files were then processed using the LD-FPG framework’s preprocessing scripts to generate the inputs for our models, including a condensed.json file that provides consistent, zero-based atom indexing and defines the atom quadruplets required for calculating all backbone and side-chain dihedral angles.
Alanine Dipeptide.
This dataset features N-acetyl-L-alanine-N’-methylamide, a 22-atom molecule commonly known as alanine dipeptide. It is a canonical benchmark for developing and testing new simulation methods due to its simple yet non-trivial conformational landscape, which is primarily described by its two backbone dihedral angles ($\phi,\psi$). The data was sourced from the CMB data repository at ftp.imp.fu-berlin.de and consists of a 250 ns simulation trajectory with solvent molecules removed.
Adenosine A1 Receptor (A1AR).
[…]
Adenosine A2A Receptor (A2AR).
Our second GPCR test case was the human adenosine A2A receptor, another Class A GPCR that serves as a key model system for studying receptor activation mechanisms. The data corresponds to a simulation of the receptor in its apo (ligand-free) state and was sourced from a Zenodo record (DOI: 10.5281/zenodo.13460724) supplementary to a detailed study on its dynamics.
S1.3 Qualitative rollouts: dihedral flips in a dipeptide and TM6 motion in a GPCR
To complement the quantitative metrics, Fig. S3 shows structural snapshots taken directly from our latent-space rollouts. The top panel illustrates alanine–dipeptide; the bottom panel shows the A1AR GPCR. These images link the latent dynamics to familiar structural changes: local backbone dihedral flips in a small molecule and the hallmark outward displacement of transmembrane helix 6 (TM6) in a receptor.
Alanine–dipeptide (Koopman).
The upper snapshots are sampled at evenly spaced steps from a Koopman rollout. Across frames, the molecule visits distinct conformers driven by rotations about the backbone dihedrals $(\phi,\psi)$, covering multiple metastable Ramachandran regions rather than collapsing to a single geometry. This visual diversity is consistent with the small backbone JSD reported in Table 1 and the ensemble analysis in Sec. 4.2, while also reflecting Koopman’s tendency to slightly damp fluctuations.
A1AR (Autoregressive NN).
The lower snapshots are taken from a single long rollout of the autoregressive neural propagator, again at evenly spaced frames. The dashed circle marks the intracellular end of TM6. Over time TM6 swings outward from the receptor core while the seven-helix bundle remains well-folded, matching the activation-associated opening on the cytoplasmic side. This qualitative progression mirrors the quantitative TM3–6/TM3–7 free-energy analysis in Fig. 4, where the NN tracks the principal low-energy valley and samples the transition corridor without structural collapse.
Figure S3: Representative structural snapshots from latent rollouts. Top: Alanine–dipeptide conformers sampled from a Koopman rollout show backbone dihedral changes across frames. Bottom: A1AR snapshots from an autoregressive NN rollout; the dashed circle highlights the intracellular end of TM6, which moves outward over time—a hallmark of GPCR activation. These qualitative views align with the ensemble statistics in Table 1 and the TM-distance thermodynamics in Fig. 4.
S1.4 Supplementary Tables
Table S1: Summary of notation used in the description of the propagator models.
Symbol Description
$z_{t}\in\mathbb{R}^{d}$ Latent space vector at time $t$.
$\mathbf{X},\mathbf{Y}\in\mathbb{R}^{d\times(M-1)}$ Snapshot matrices with columns as time: $\mathbf{X}=[z_{0},\dots,z_{M-2}]$, $\mathbf{Y}=[z_{1},\dots,z_{M-1}]$.
$A\in\mathbb{R}^{d\times d}$ Koopman linear operator.
$f_{\theta}$ Autoregressive neural network with parameters $\theta$.
$p(z)$ Equilibrium probability distribution of the latent variable $z$.
$s(z)=\nabla_{z}\log p(z)$ Score function of the equilibrium distribution.
$\epsilon_{\theta}(z_{t},\tau)$ Pre-trained LD-FPG diffusion model (denoiser).
$\sigma_{\tau}$ Noise schedule standard deviation from the diffusion model at step $\tau$.
$\Delta t$ Integration time step for Langevin dynamics.
$T$ Temperature parameter for Langevin dynamics.
$\eta_{t}$ Stochastic noise term, typically $\eta_{t}\sim\mathcal{N}(0,\sigma^{2}I)$.
Table S2: Glossary of Biophysical Terms. Definitions of key concepts from molecular dynamics and structural biology used in this work, tailored for a machine learning audience.
Term Description for an ML Audience
Collective Variable (CV) A low-dimensional function of atomic coordinates (e.g., a distance or angle) designed to capture a specific, slow dynamic process like protein folding. Traditional simulation methods often require pre-defining good CVs; our work uses a learned latent space to discover them automatically.
Potential of Mean Force (PMF) Essentially, an effective "energy landscape" for a molecule in solution. Lower values correspond to more probable (stable) conformations. It is the target distribution that score-guided Langevin dynamics aims to sample from, often visualized as a "free-energy surface".
Dihedral Angles ($\phi,\psi,\chi$) Rotational angles around covalent bonds that define the geometry of a molecule. $\phi$ (phi), $\psi$ (psi): Define the rotation of the protein backbone. $\chi$ (chi): Define the rotation of the amino acid side-chains. Their statistical distributions are a sensitive measure of structural fidelity.
Ramachandran Plot A 2D plot of the backbone dihedral angles ($\phi,\psi$). Certain regions of this plot are "allowed" based on steric constraints, leading to characteristic high-probability basins that correspond to stable secondary structures like alpha-helices and beta-sheets.
Rotamer A discrete, low-energy, and therefore highly probable conformation of a protein’s side-chain, defined by its set of $\chi$ angles. A key test for generative models is whether they can reproduce the correct statistical distribution of these rotameric states.
GPCR G protein-coupled receptor. A large and important family of transmembrane proteins that act as cellular signal transducers. They are highly dynamic and switch between different functional states (e.g., inactive, active), making them an ideal and challenging test system for dynamic models.
lDDT local Distance Difference Test. A metric for assessing the quality of a protein structure prediction by evaluating how well local inter-atomic distances are preserved relative to a reference structure. Unlike RMSD, it is less sensitive to global rotations and more focused on local geometric accuracy.
RMSF Root-Mean-Square Fluctuation. For each atom, this metric calculates the standard deviation of its position over time in a simulation trajectory. It measures the "amplitude of motion" or flexibility of different parts of the protein.